# Equivariant Spherical Transformer for Efficient Molecular Modeling

### v2

@@ 3.2.2 Spherical Attention
Treating $\mathbf{f}^{*}$ as a sequential data, where each sampled point $s$ has a $C$-dimensional feature vector, we apply a attention mechanism vaswani2017attention :
@@
where the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices are obtained through point-wise linear (PW-Linear) transformations of $\mathbf{f}^{*}$. The spherical attention (SA) is the core of EST. We also incorporate the pre-LayerNorm strategy xiong2020layer and feedforward networks (FFN) in EST. Our EST is applied in two ways: (i) for interactions between two spatial representations, $\mathbf{Q}$ is derived from one, and $\mathbf{K}$ and $\mathbf{V}$ from the other; (ii) for learning features within a single spatial representation, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from it.
@@
Assuming no information loss during the spherical Fourier transform and its inverse, i.e., $\mathbf{x}=\mathcal{F}^{-1}\mathcal{F}(\mathbf{x})$, Theorem 1 can evaluate the expressive power of EST.
Theorem 1.
For any two steerable representations $\mathbf{u}\in\mathbb{V}_{0\to l_{1}}$ and $\mathbf{v}\in\mathbb{V}_{0\to l_{2}}$, the spatial representation of their CG tensor product, $\mathbf{u}\otimes\mathbf{v}$, can be uniformly approximated by EST modules operating on the spatial representations of $\mathbf{u}$ and $\mathbf{v}$.
The proof can be found in Appendix B.1. We note that a Fourier transform without information loss is attainable by satisfying the Nyquist sampling rate, which mandates a minimum of $(2L)^{2}$ sampling points on the sphere zitnick2022spherical . Therefore, Theorem 1 substantiates the potential for substituting tensor products with EST. Furthermore, we present an empirical comparison of the expressiveness of EST and tensor product-based operations in Section 4.3, where our EST significantly surpasses the theoretical upper bound of tensor products.
Equivariance
We demonstrate that the equivariance of EST is related to the uniformity of the sampling points in Fourier transform.
Theorem 2.
If the sampling point density is the same across any local region of the sphere, EST operation preserves SO(3)-equivariance:
$\mathcal{F}^{-1}\big{(}\mathrm{EST}\big{(}\mathcal{F}(\mathbf{Dx})\big{)}\big{)}=\mathbf{D}\mathcal{F}^{-1}\big{(}\mathrm{EST}\big{(}\mathcal{F}(\mathbf{x})\big{)}\big{)},$ (10)
where $\mathbf{D}$ represents an arbitrary Wigner-D rotation matrix.
The proof is provided in Appendix B.2. Most prior works passaro2023reducing ; liao2024equiformerv implement spherical Fourier transforms using the e3nn library e3nn . However, e3nn does not guarantee a uniform sampling point density across all regions of the sphere. To address this, we redefine the sampling implementation using Fibonacci Lattices (FL) gonzalez2010measurement . The Cartesian coordinate of each sampling point $s$ is defined by
where $z_{s}=\frac{2s-1}{S-1}$ and $\lambda=\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. The Fourier basis functions must satisfy orthogonality, i.e., $\int_{\Omega}Y^{(l,m)}(\vec{\mathbf{p}})Y^{(l^{\prime},m^{\prime})*}(\vec{\mathbf{p}})d\Omega=\delta_{ll^{\prime}}\delta_{mm^{\prime}}$, where $\delta_{ij}$ is the Kronecker delta. Consequently, we redefine the implementation of conjugate functions for the inverse transform as:
where $\lambda^{(l,m)}=1/\sum_{s}|Y^{(l,m)}(\vec{\mathbf{p}}_{s})|^{2}$. We provide detailed discussions between our sampling and the e3nn implementation in the Appendix B.2. In addition, we conducted ablation experiments in Section 4.3 to demonstrate that sampling methods with poor uniformity can severely compromise equivariance.
Relative Orientation Embedding on the Sphere
Despite its strong expressiveness, EST lacks the inherent ability to capture orientation relationships between sampling points, potentially leading to ambiguities dufter2022position . For example, if the features at points $s_{1}$ and $s_{2}$ are identical ($\mathbf{f}^{*}_{s_{1}}=\mathbf{f}^{*}_{s_{2}}$), their contributions to other sampling points $s_{i}$ become indistinguishable. To address this, we introduce a relative orientation embedding by augmenting the query $\mathbf{Q}$ and key $\mathbf{K}$ vectors with orientation information:
@@
Consequently, the inner product between $\mathbf{Q}_{s_{i}}$ and $\mathbf{K}_{s_{j}}$ now incorporates a term aware of the orientation, related to $\vec{\mathbf{p}}^{T}_{s_{i}}\vec{\mathbf{p}}_{s_{j}}$. Moreover, this augmentation does not compromise the equivariant inductive bias of EST, as proven in Appendix B.3. The intuition behind our approach shares similarity with rotary position embeddings in NLP su2024roformer , where relative position embedding should be invariant under transformation. However, while NLP focuses on permutation invariance, our approach addresses rotation invariance, necessitating a fundamentally different implementation.
Figure 2: The architecture and building blocks of EST. SH and DTP denote spherical harmonic embedding and depth-wise tensor product liao2023equiformer , respectively. For simplicity, the Fourier and inverse Fourier transform steps are omitted. (a) Overall architecture. (b) Message block. (c) Update block. (d) Two experts operating on the steerable and spatial representations, respectively.
@@ 3.2.3 Mixture of Hybrid Experts
To enhance the model’s representational capacity, we utilize equivariant FFNs within a Mixture of Experts (MoE) framework shazeer2017outrageously . Starting with the steerable representation $\mathbf{x}$, the invariant ($l=0$) component is used to compute expert weights through a gate function:
@@
where $\mathbf{W}_{G}\in\mathbb{R}^{C\times E}$ is a trainable weight matrix. As illustrated in Figure 2(c), we employ two distinct expert structures: (1) a steerable FFN, operating on steerable inputs and consisting of two DW-Linear layers with an intermediate Gate activation liao2023equiformer ; and (2) a spherical FFN, processing spatial inputs and comprising two PW-Linear layers with an intermediate SiLU activation elfwing2018sigmoid . The final output is a weighted combination of the outputs from these hybrid experts, with the weights determined by Equation (14). These two expert types offer a crucial trade-off: the steerable FFN guarantees strict equivariance, while the spherical FFN provides enhanced expressive power. We therefore adapt their numbers based on the specific demands of the task. Additionally, we investigate common MoE techniques, including balancing loss lepikhin2020gshard and shared experts dai2024deepseekmoe . Further details and ablation studies can be found in Appendix C.
3.3 Overall Architecture
EST can be employed both in the message block to compute edge features and in the update block to refine node embeddings. Figure 2(b,c) defines a message passing layer with EST. When computing the message, the $\mathbf{Q}$ in SA come from the aggregation of node $i$ and $j$ embeddings, while $\mathbf{K}$ and $\mathbf{V}$ is derived from the spherical harmonic representation of the relative position $\vec{\mathbf{r}}_{ij}$. In the SA of the update block, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from the aggregated message.
EST serves as a flexible building block and can also be integrated into other equivariant models. Since the primary function of EST is to extract critical interactions or geometric features from complex high-degree representations, we prefer to replace the update blocks in existing models with our EST-based update block, where the input (aggregated message) contains sufficient and minimally processed neighbor features. With this strategy, EST-based architectures can also benefit from the computational efficiency of lightweight message blocks, while keep advantages to enhance the expressiveness.
@@ 4 Experiments
In this section, we construct experiments to investigate the efficacy of the proposed method. We evaluate its performance on the S2EF/IS2RE tasks from the OC20 chanussot2021open benchmark and QM9 ramakrishnan2014quantum tasks. Our analysis involves a comparison with a wide range of baseline models, including Schnet schutt2018schnet , PaiNN schutt2021equivariant , SEGNN brandstetter2021geometric , TFN thomas2018tensor , Dimenet++ klicpera2020fast , Equiformer liao2023equiformer , and Equiformerv2 liao2024equiformerv . Specifically for the OC20 benchmark, we extend our comparison to include SpinConv shuaibi2021rotation , GemNet gasteiger2021gemnet ; gasteiger2022gemnetoc , SphereNet liu2021spherical , SCN zitnick2022spherical and eSCN passaro2023reducing . Similarly, our QM9 experiments are augmented with comparisons to TorchMD-NET tholke2022equivariant , EQGAT le2022equivariant and GotenNet aykentgotennet . The specific configurations employed for each baseline model can be found in Appendix D.1.
Table 1: S2EF results of models trained on S2EF-All training dataset. $\lambda_{E}$ is the coefficient of the energy loss. “All+MD” denotes training models with additional OC20 MD dataset.
@@
SchNet 9.1M - $548.525$ $56.8$
DimeNet++-L-F+E 10.7M 4.6 $515.475$ $32.75$
SpinConv 8.5M 6.0 371.011 65371.011\,65 41.197 87541.197\,875
GemNet-dT 32M 25.8 315.118 775315.118\,775 27.172 47527.172\,475
GemNet-OC 39M 18.3 244.289 799 999 999 99244.289\,799\,999\,999\,99 21.719 050 000 000 00321.719\,050\,000\,000\,003
SCN (20 layers, test) 271M - $244$ $17.7$
eSCN (20 layers, test) 200M 2.9 $242$ $17.1$
@@
EquiformerV2 ($\lambda_{E}=2$, 20 layers) 153M 1.8 $236$ 15.7
EST ($\lambda_{E}=4$, 8 layers) 45M 6.8 231 $16.1$
@@ 4.1 OC20 Results
The OC20 dataset, comprising over 130 million molecular structures with force and energy labels, covers a broad spectrum of materials, surfaces, and adsorbates. We evaluate our models on two core sub-datasets of OC20: S2EF and IS2RE. All experimental configurations are detailed in Appendix D.2. Notably, We did not use very deep models or very long training schedules. For instance, our S2EF experiments utilize only 8 layers, and our IS2RE experiments employ 6 layers, significantly fewer than many advanced methods. Furthermore, as mentioned in Section 3.3, to enhance computational efficiency and facilitate a more direct comparison of different fundamental modules, we replace the EST architecture’s message module with the graph attention modules from EquiformerV2 and Equiformer for S2EF and IS2RE tasks, respectively.
@@
We trained an 8-layer EST model on the S2EF-All dataset and evaluated its performance on the S2EF validation sets. Each validation set was divided into four similarly-sized subsets: ID, OOD Ads, OOD Cat, and OOD Both. Consistent with prior work passaro2023reducing ; liao2024equiformerv , the best model was selected based on its performance on the ID validation subset during training and subsequently evaluated across all validation subsets. For comparison with SCN and eSCN, we report their test results due to the absence of publicly available validation results; previous work liao2024equiformerv indicates that the validation and test sets have a similar distribution. We compared our results with two variants of the SOTA EquiformerV2 model: a deep 20-layer architecture and one trained with additional MD dataset. As shown in Table 1, EST outperformed both EquiformerV2 variants in energy prediction. Furthermore, EST achieved competitive results in force prediction, surpassing several deeper baselines including SCN and eSCN.
@@ 4.2 QM9 Results
Given that the QM9 dataset is considerably smaller than OC20, models are more prone to overfitting if equivariance is destroied. Ablation study in Table 5 shows that practical implementation for Fourier transform may lead to minor deviations from perfect equivariance. Nevertheless, EST and EST (with GA) achieve best on eight of the twelve tasks (see Table 3). Notably, EST (with GA) significantly outperformed Equiformer despite utilizing a similar message block. Furthermore, we included a comparison with GotenNet aykent2025gotennet , which employs a more extensive training schedule. To ensure a fair assessment, we applied this identical schedule to EST as well; see Appendix D.3 for details.
@@ 4.3 Ablation study
Following joshi2023expressive , we employ evaluation metrics that distinguish $n$-fold symmetric structures to precisely assess EST’s expressive power. We first construct a single-layer message-passing layer based on Figure 2, consistent with tensor product-based operations (TFN, MACE). Additionally, we remove the steerable FFN for a clear comparison. As demonstrated in dym2021on ; joshi2023expressive , the expressive power of tensor product-based models is limited by the maximum degree $L$, failing to perfectly distinguish $n$-fold symmetric structures when $n>L$. From Table 4, we observe that EST significantly breaks through this theoretical boundary. It allows the model to distinguish symmetric structures even with very high fold (e.g., using 1-degree EST for 100-fold structures).
@@
While we theoretically demonstrate that EST can achieve strict equivariance, perfect uniform sampling is challenging in practice, potentially leading to loss of equivariance. To investigate it, we build untrained networks with 1 to 6 layers based on Figure 2 and control the number of spherical sampling points in the Fourier transforms within the message and update blocks. We randomly select 1000 molecules from QM9, compute their outputs $y_{1},...,y_{1000}$, and then compute the outputs after applying random rotations, $\hat{y}_{1},...,\hat{y}_{1000}$. The average absolute error $(1/1000)\sum_{i=1}^{1000}|y_{i}-\hat{y}_{i}|$ serves as our measure of equivariance loss. As shown in Table 5, EST with FL sampling closely approximates strict equivariance even without training, and its equivariance further improves with an increasing number of spherical sampling points. Additionally, EST with the sampling implementation in e3nn show a poor equivariance.

### v3

@@ 3.2.2 Spherical Attention
To prepare the data, we flatten all sampled points from $\mathbf{f}^{*}$ into a sequential format, where each point $s$ represented by a $C$-dimensional feature vector. We then apply an attention mechanism [26]:
@@
where the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices are obtained through point-wise linear (PW-Linear) transformations of $\mathbf{f}^{*}_{s_{i}}$. The spherical attention is the core of EST. We also incorporate the pre-LayerNorm strategy [27] and feedforward networks (FFN) in EST. Our EST is applied in two ways: (i) for interactions between two spatial representations, $\mathbf{Q}$ is derived from one, and $\mathbf{K}$ and $\mathbf{V}$ from the other; (ii) for learning features within a single spatial representation, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from it.
Equivariance and sampling strategies
We demonstrate that the equivariance of the EST module is contingent upon the uniformity of the sampling points used in the FT. We begin with the following theorem:
Theorem 1.
Let $\mathcal{P}=\{\vec{\mathbf{p}}_{1},\dots,\vec{\mathbf{p}}_{N}\}$ be a set of sampled points used for the FT. The EST module is strictly SO(3)-equivariant if the set $\mathcal{P}$ is closed under arbitrary rotations from the SO(3) group. That is, for any point $\vec{\mathbf{p}}_{i}\in\mathcal{P}$ and any rotation $\mathbf{R}\in\text{SO}(3)$, there exists a point $\vec{\mathbf{p}}_{j}\in\mathcal{P}$ such that $\vec{\mathbf{p}}_{j}=\mathbf{R}\vec{\mathbf{p}}_{i}$. In this case, the equivariance of EST property holds:
$\mathcal{F}^{-1}\big(\mathrm{EST}\big(\mathcal{F}(\mathbf{D}\mathbf{x})\big)\big)=\mathbf{D}\mathcal{F}^{-1}\big(\mathrm{EST}\big(\mathcal{F}(\mathbf{x})\big)\big),$ (10)
where $\mathbf{D}$ is an arbitrary Wigner-D rotation matrix.
The proof is provided in Appendix B.1. Due to the inherent symmetry of the sphere, a uniform sampling strategy can approximate the closed property of the point set $\mathcal{P}$, meaning a subset of rotations will satisfy $\vec{\mathbf{p}}_{j}=\mathbf{R}\vec{\mathbf{p}}_{i}$. In contrast, most prior works [15, 16] implement spherical FT using the e3nn library [17], which results in point densities that vary drastically with the polar angle $\theta$ (see Figure 2(a) for visualization). To address this, we redefine the sampling implementation using Fibonacci Lattices (FL) [28]. The Cartesian coordinate of each sampling point $s$ is defined by
where $z_{s}=\frac{2s-1}{S-1}$ and $\lambda=\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. The (11) ensures the sampled points are distributed on the spherical surface, and different regions have similar densities (see Figure 2(b)). While FL sampling provides a high degree of uniformity, we further enhance this by simulating the points’ motion on the sphere through consistent repulsions. This process, which resembles a simple molecular dynamics simulation, where each atom-like point is confined to the spherical surface, with forces determined by the distances between them.
The orthogonality of the basis functions in (4) is defined by $\int_{\Omega}Y^{(l,m)}(\vec{\mathbf{p}})Y^{(l^{\prime},m^{\prime})*}(\vec{\mathbf{p}})d\Omega=\delta_{ll^{\prime}}\delta_{mm^{\prime}}$, where $\delta_{ij}$ is the Kronecker delta. Consequently, we redefine the implementation of the conjugate functions for the inverse FT as:
where $\lambda^{(l,m)}=1/\sum_{s}|Y^{(l,m)}(\vec{\mathbf{p}}_{s})|^{2}$. A detailed discussion comparing our sampling method with the e3nn implementation is provided in Appendix B.1. Furthermore, our ablation experiments (Table 5) demonstrate that sampling methods with poor uniformity can severely compromise equivariance.
@@
As demonstrated by [11, 12, 13], the expressive power of CG tensor product-based frameworks is constrained by the maximum degree $L$, a finding corroborated by ablation experiments (Table 4). In contrast, our proposed EST module can project steerable representations of any degree onto the sphere and perform flexible, learnable operations. This allows the resulting features to be reprojected back to any desired degree, thereby potentially overcoming the expressiveness limitations of tensor products.
Proposition 2.
Under the assumption of no information loss in the spherical Fourier transform and its inverse, the function space spanned by the spherical Transformer encompasses that spanned by the CG tensor product.
The detail is provided in Appendix B.2. We note that a spherical FT with no information loss is attainable by satisfying the Nyquist sampling rate, which mandates a minimum of $(2L)^{2}$ sampling points on the sphere [14]. Thus, Proposition 2 validates the potential for substituting tensor products with EST. An empirical comparison of the expressiveness of EST and tensor product-based operations is presented in Section 4.3, where our EST significantly surpasses the upper bound of tensor products.
Relative Orientation Embedding
In its basic form, spherical attention determines the dependencies between points based solely on their feature values, ignoring their relative positions in the input sequence. This means the output feature of a given point remains unchanged even if the order of all other points is altered. Such behavior is undesirable, as different point orderings correspond to distinct spherical functions. To enable the model to distinguish between these configurations, we introduce a relative orientation embedding by augmenting the query ($\mathbf{Q}$) and key ($\mathbf{K}$) vectors with orientation information:
@@
Consequently, the inner product between $\mathbf{Q}_{s_{i}}$ and $\mathbf{K}_{s_{j}}$ now incorporates a term related to the orientation-aware inner product $\vec{\mathbf{p}}^{T}_{s_{i}}\vec{\mathbf{p}}_{s_{j}}$. Crucially, this augmentation does not compromise the equivariant inductive bias of EST, as proven in Appendix B.3. Our approach shares conceptual similarities with Rotary Position Embeddings (RoPE) in NLP [29], where the relative position embedding is invariant under transformation. However, our method fundamentally differs from RoPE by addressing rotation invariance instead of permutation invariance.
@@ 3.2.3 Mixture of Hybrid Experts
To enhance the model’s representational capacity, the FFNs in EST is reformed within a Mixture of Experts (MoE) framework [30]. Starting with the steerable representation $\mathbf{x}$, the invariant ($l=0$) component is used to compute expert weights through a gate function:
@@
where $\mathbf{W}_{G}\in\mathbb{R}^{C\times E}$ is a trainable weight matrix. We then employ two distinct expert structures: (1) a steerable FFN, operating on steerable inputs and consisting of two DW-Linear layers with an intermediate Gate activation [10]; and (2) a spherical FFN, processing spatial inputs and comprising two PW-Linear layers with an intermediate SiLU activation [31]. The final output is a weighted combination of the outputs from these hybrid experts, with the weights determined by (14). These two expert types offer a crucial trade-off: the steerable FFN guarantees strict equivariance, while the spherical FFN provides enhanced expressive power. We therefore adapt their numbers based on the specific demands of the task. Further details and ablation studies can be found in Appendix C.
3.3 Applying the EST Module to Equivariant Models
We can apply EST module into equivariant models by two primary ways:
  • Message block: We replace the conventional message block, often a tensor product like $\mathrm{Combine}(\mathbf{x}_{i},\mathbf{x}_{j})\otimes\vec{\mathbf{r}}_{ij}$, with a EST. Here, the query ($\mathbf{Q}$) is formed from a combination of the embeddings of nodes $i$ and $j$, while the key ($\mathbf{K}$) and value ($\mathbf{V}$) are derived from the spherical harmonic representation of the relative position vector $\vec{\mathbf{r}}_{ij}$.
  • Update block: To replace update block, all three attention components—query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$)—are derived directly from the aggregated messages.
We provide an architecture in Appendix C.4 that uses the strategies mentioned above. Moreover, our approach leverages EST as a plug-and-play module that integrates seamlessly into existing equivariant models. Specifically, we recommend replacing the standard update blocks—which require strong expressiveness to learn richer representations from complex aggregated message—with our EST-based update blocks. This allows EST-based architectures to benefit from the computational efficiency and stability of existing message blocks while still enhancing their expressive power.
@@ 4 Experiments
In this section, we construct experiments to investigate the efficacy of the proposed method. We evaluate its performance on the S2EF/IS2RE tasks from the OC20 [18] benchmark and QM9 [19] tasks. Our analysis involves a comparison with a wide range of baseline models, including Schnet [5], PaiNN [32], SEGNN [4], TFN [3], Dimenet++ [33], Equiformer [10], Equiformerv2 [16] and EScAIP [34]. Specifically for the OC20 benchmark, we extend our comparison to include SpinConv [35], GemNet [9, 36], SphereNet [37], SCN [14] and eSCN [15]. Similarly, our QM9 experiments are augmented with comparisons to TorchMD-NET [38], EQGAT [39] and GotenNet [23]. The specific configurations employed for each baseline model can be found in Appendix D.1.
Table 1: S2EF validation results. $\lambda_{E}$ is the coefficient of the energy loss. “only All” denotes training models with only S2EF-All dataset, and “All+MD” introduces additional OC20 MD dataset for training.
@@
SchNet (only All) 9.1M - $548.525$ $56.8$
DimeNet++-L-F+E (only All) 10.7M 4.6 $515.475$ $32.75$
SpinConv (only All) 8.5M 6.0 371.011 65371.011\,65 41.197 87541.197\,875
GemNet-dT (only All) 32M 25.8 315.118 775315.118\,775 27.172 47527.172\,475
GemNet-OC (only All) 39M 18.3 244.289 799 999 999 99244.289\,799\,999\,999\,99 21.719 050 000 000 00321.719\,050\,000\,000\,003
SCN (20 layers, only All, test results) 271M - $244$ $17.7$
eSCN (20 layers, only All, test results) 200M 2.9 $242$ $17.1$
EScAIP-Small (All+MD) 83M - 229 $15.1$
@@
EquiformerV2 ($\lambda_{E}=2$, 20 layers, only All) 153M 1.8 $236$ $15.7$
EST ($\lambda_{E}=4$, 8 layers, only All) 45M 6.8 229 $15.8$
EST ($\lambda_{E}=2$, 8 layers, only All) 45M 6.8 $232$ 15.0
@@ 4.1 OC20 Results
Dataset and Configurations
The OC20 dataset, comprising over 130 million molecular structures with force and energy labels, covers a broad spectrum of materials, surfaces, and adsorbates. We evaluate our models on two core sub-datasets of OC20: S2EF and IS2RE. All experimental configurations are detailed in Appendix D.2. Notably, We did not use very deep models or very long training schedules. For instance, our S2EF experiments utilize a 8-layer architecture, and IS2RE experiments employ a 6-layer one, significantly fewer than many advanced methods. Furthermore, as mentioned in Section 3.3, to enhance computational efficiency and facilitate a more direct comparison of different fundamental modules, we replace the EST architecture’s message module with the graph attention modules from EquiformerV2 and Equiformer for S2EF and IS2RE tasks, respectively.
@@
We trained an 8-layer EST model on the S2EF-All dataset and evaluated its performance on the S2EF validation sets. Each validation set was divided into four similarly-sized subsets: ID, OOD Ads, OOD Cat, and OOD Both. Consistent with prior work [15, 16], the best model was selected based on its performance on the ID validation subset during training and subsequently evaluated across all validation subsets. For comparison with SCN and eSCN, we report their test results due to the absence of publicly available validation results; previous work [16] indicates that the validation and test sets have a similar distribution. We compared our results with two variants of the EquiformerV2 model: a deep 20-layer architecture and one trained with additional MD dataset. As shown in Table 1, EST outperformed both EquiformerV2 variants in energy and forces prediction. Furthermore, EST achieved competitive results in force prediction, surpassing several bigger baselines including EScAIP, SCN and eSCN.
@@ 4.2 QM9 Results
Results
Given that the QM9 dataset is considerably smaller than OC20, models are more prone to overfitting if equivariance is destroied. Ablation study in Table 5 shows that practical implementation for Fourier transform may lead to minor deviations from perfect equivariance. Nevertheless, EST and EST (with GA) achieve best on eight of the twelve tasks (see Table 3). Notably, EST (with GA) significantly outperformed Equiformer despite utilizing a similar message block. Furthermore, we included a comparison with GotenNet [25], which employs a more extensive training schedule. To ensure a fair assessment, we applied this schedule to EST as well; see Appendix D.3 for details.
@@ 4.3 Ablation study
Following [12], we employ evaluation metrics that distinguish $n$-fold symmetric structures to precisely assess EST’s expressive power. We first construct a single-layer message-passing layer based on Figure 3, consistent with tensor product-based operations (TFN, MACE). Additionally, we remove the steerable FFN for a clear comparison. As demonstrated in [11, 12], the expressive power of tensor product-based models is limited by the maximum degree $L$, failing to perfectly distinguish $n$-fold symmetric structures when $n>L$. From Table 4, we observe that EST significantly breaks through this theoretical boundary. It allows the model to distinguish symmetric structures even with very high fold (e.g., using 1-degree EST for 100-fold structures).
@@
While we theoretically demonstrate that EST can achieve strict equivariance, perfect uniform sampling is challenging in practice, potentially leading to loss of equivariance. To investigate it, we build untrained networks with 1 to 6 layers based on a EST-only message passing, controlling the number of spherical sampling points and choosing whether to use molecular dynamics-like optimization. We randomly select 1000 molecules from QM9, compute their outputs $y_{1},...,y_{1000}$, and then compute the outputs after applying random rotations, $\hat{y}_{1},...,\hat{y}_{1000}$. The average absolute error $(1/1000)\sum_{i=1}^{1000}|y_{i}-\hat{y}_{i}|$ serves as our measure. As shown in Table 5, EST with FL sampling closely approximates strict equivariance even without training, and its equivariance further improves with an increasing number of sampling points or a dynamics-like optimization. Additionally, we evaluated the equivariance error of 6-layer trained EST models on 1000 data samples from the each task of QM9 dataset. Across all properties, the average error was $1.8\times 10^{-5}\pm 0.7\times 10^{-5}$.